# Monitored long-range interacting systems: spin-wave theory for quantum trajectories

### v1

@@ III.2.1 Infinitesimal increments
The stochastic master equation (1) presented in Sec. II determines the time evolution of the expectation value of operators. For a time-independent operator $\hat{O}$, the expectation $\langle\hat{O}\rangle={\rm{Tr}}[\hat{\rho}\hat{O}]$ evolves as follows,
@@
In the equations above, the Hamiltonian $\hat{H}$ and the dissipation operators $\hat{L}_{i}$ should be expressed in terms of the bosonic operators $\hat{b}_{i}$ using the substitution rules defined in Eqs. (19) and (21). The Gaussian approximation then allows evaluating the expectation value of every term beyond quadratic order in $\hat{b}_{i}$ in terms of one- and two-point correlators, i.e. $\beta_{i}$, $u_{ij}$ and $v_{ij}$, thanks to the Wick theorem. The increments $\delta\beta_{i}$, $\delta u_{ij}$, and $\delta v_{ij}$ are then calculated using the discretized versions of the equations above. This can be achieved with the substitution ${\rm{d}}t\to\delta t$ for a sufficiently small time step $\delta t$. The noise is approximated with ${\rm{d}}\mbox{w}_{i}\to\delta\mbox{w}_{i}=\sqrt{\delta t}(X_{i}+{\rm{i}}Y_{i})$, where the pair of random vectors $(\mathbf{X},\mathbf{Y})$ at every time step is drawn from a multivariate real Gaussian distribution with zero mean and the following covariance matrix  [^1] Note that it can be computationally more practical to diagonalize the dissipator first (see Appendix A) and work in the basis where the noise becomes diagonal, in which case the covariance matrix $\mathbf{K}$ also diagonalizes and it suffices to generate independent Gaussian random numbers to approximate the noise,
@@ III.2.2 Re-alignment of the frame
as a direct implication of Eqs. (20) and (21). This can be achieved by moving the frame smoothly along a path $\theta(\tau),\phi(\tau)$ parametrized by some parameter $\tau$. The unitary transformation defined in Eq. (16) therefore becomes $\tau$-dependent,
@@
This equation can be integrated analytically by considering the path on the Bloch sphere from $(\theta_{1},\phi_{1})$ to $(\theta_{2},\phi_{2})$ following the two segments [as illustrated in 1 (b)]:
@@
In the first segment $\gamma_{1}$, we are following a latitude line of constant $\theta$, and the evolution equation simplifies to be a separable first-order differential equation of $\hat{b}_{i}(\phi)_{\theta=\theta_{1}}$:
${\rm{d}}\hat{b}_{i}(\phi)={\rm{i}}\sqrt{\dfrac{s}{2}}\sin\theta_{1}{\rm{d}}\phi-{\rm{i}}\cos\theta_{1}\hat{b}_{i}{\rm{d}}\phi\,,$ (37)
which can be solved as
$\hat{b}_{i}(\theta_{1},\phi_{2})=$ $\leavevmode\nobreak\ \mathopen{}\mathclose{{}\left[\sqrt{\dfrac{s}{2}}\tan\theta_{1}+\hat{b}_{i}(\theta_{1},\phi_{1})}\right]{\rm{e}}^{-{\rm{i}}\Delta\phi\cos\theta_{1}}$ (38)
$-\sqrt{\dfrac{s}{2}}\tan\theta_{1}\,,$
where $\Delta\phi\equiv\phi_{2}-\phi_{1}$. The evolution along the second segment, following a longitude line of constant $\phi$, is simply
${\rm{d}}\hat{b}_{i}(\theta)=-\sqrt{\dfrac{s}{2}}{\rm{d}}\theta\,,$ (39)
which gives
$\hat{b}_{i}(\theta_{2},\phi_{2})$ $=\hat{b}_{i}(\theta_{1},\phi_{2})-\sqrt{\dfrac{s}{2}}\Delta\theta$ (40)
$=$ $\leavevmode\nobreak\ \mathopen{}\mathclose{{}\left[\sqrt{\dfrac{s}{2}}\tan\theta_{1}+\hat{b}_{i}(\theta_{1},\phi_{1})}\right]{\rm{e}}^{-{\rm{i}}\Delta\phi\cos\theta_{1}}$
@@
where $\Delta\theta\equiv\theta_{2}-\theta_{1}$. This equation immediately implies the evolution of $\underline{\beta}$, upon taking the expectation of both sides,
@@
Note that these increments are not necessarily infinitesimal due to the coordinate singularities near the poles of the Bloch sphere, where the angle $\phi$ may change drastically even for physically small movements of the frame. With the angular increments in hand, we are now ready to update the remaining parameters, namely the one- and two-point correlators for the bosonic Gaussian state. The first moments $\beta_{i}$ can be updated directly using Eq. (40) with the operator $\hat{b}_{i}$ replaced by its expectation value $\beta_{i}$. This equation also implies the evolution of the fluctuation operators $\hat{\delta}_{i}=\hat{b}_{i}-\beta_{i}$, which is simply
@@
Lastly, note that our method presented above is fully generic, as it suffices to systematically apply the bosonization and the Gaussian approximation to obtain the equations of motion. Therefore, it can be straightforwardly extended to other types of systems beyond those we considered. An example of its generalization to spin-boson systems is presented in Appendix D together with some illustrative numerical results.
IV Results
In this section, we apply the presented theory of spin-wave quantum trajectories to study the power-law interacting spin model introduced in Sec. II.1. We focus on $s=1/2$ and fix the interaction strength at $J=0.1\kappa$ for the rest of the section.
IV.1 Infinite-range case: $\alpha=0$
We now test the quality of the approximation by evaluating both linear (e.g., observables) and non-linear (e.g., entanglement entropy) functions of the state. In order to benchmark the method, we first revisit the known case of $\alpha=0$, where the $z-z$ interaction is all-to-all with no spatial resolution. In this regime, the system can be effectively represented as a single spin $S=N/2$, and the Hamiltonian (7) becomes
$\hat{H}=\omega\hat{S}^{x}+\dfrac{2J}{S}(\hat{S}^{z})^{2}\,.$ (46)
This allows us to benchmark the spin-wave method against the exact simulation of the dynamics of the collective mode in the Dicke basis. Note that the computational complexity of our spin-wave method also significantly reduces in this case since it suffices to apply the formalism in Sec. III to an effective single-body problem with spin number $S$. This boils down to calculating three increments at each step: $\delta\beta$, $\delta u$, and $\delta v$; the re-alignment step then gives the angular increments $\Delta\theta$ and $\Delta\phi$, resets $\beta$ to zero and applies a phase factor to $u$. The computation cost is therefore constant for the spin-wave method, which is still significantly more efficient than the exact simulation which requires a Hilbert space growing linearly with $N$.
Figure 2: Single-trajectory benchmark of the spin-wave method against the exact integration of the stochastic master equation on the collective model ($\alpha=0$). (a) The collective $z-$magnetization as a function of time along a single trajectory obtained from the spin-wave method (solid line) and the exact solution (dashed line). (b) The half-chain entanglement entropy along the same trajectory from the spin-wave method (solid line) and the exact solution (dashed line). Parameters: $\omega=1.25\kappa$, $J=0.1\kappa$, $S=64$, $\kappa\delta t=10^{-4}$.
Figure 3: Benchmark of the spin-wave method on trajectory-averaged quantities. (a) Expectation of the collective spin vector $\hat{S}^{\alpha}$ as a function of time given by the spin-wave quantum trajectories (solid lines) and the exact solution of the master equation (14) (dashed lines). (b) Time-evolution of the averaged half-chain entanglement entropy from the spin-wave method (solid line) and the exact trajectories (dashed line). $4000$ trajectories are used for both the spin-wave method and the exact solution of the entanglement entropy. The parameters are the same as in Fig. 2.
To put the spin-wave quantum trajectories in fair comparison with the exact ones, we integrate the exact stochastic master equation (1) using the Euler-Maruyama method, and we adopt the same time step $\delta t$ and the same noise realization $\delta\mbox{w}_{i}(t)$ as used in the spin-wave calculation. Note that this is a very strict benchmark as the usual Gaussian trajectory approximation for bosonic systems typically does not match the exact trajectories, as shown in [109], and the approximation is only valid in the weak sense, i.e., for calculating ensemble averages. An example of a single-trajectory benchmark for the spin-wave method is shown in Fig. 2, where we compare trajectories for both the magnetization $\langle\hat{S}^{z}\rangle$ and the entanglement entropy  [^2] The entanglement entropy is calculated following the derivation in [7]. One can also choose any other nonlinear function of the state for this benchmark. $S_{E}$ against exact ones. The considered system has $S=64$ and $\omega=1.25\kappa$, which, as predicted by the mean-field theory, corresponds to the time-crystal phase in the thermodynamical limit. The initial state is the fully polarized Dicke state (all spins pointing up) for both simulations. (For the spin-wave method, this corresponds to initializing with $\theta=\phi=0$ and $u=v=0$.) Surprisingly, the trajectories obtained with the spin-wave method faithfully reproduce the dynamics of both quantities throughout most of the evolution. This suggests that the proposed method can be used as an approximation to resolve dynamics on the level of single trajectories.
Figure 4: Benchmark of the spin-wave method on the steady state of the collective model with $S=64$ and $J=0.1\kappa$. (a) Expectation of the collective $z$-magnetization as a function of the drive $\omega$ obtained with the spin-wave quantum trajectories and the exact solution (see legend). The mean-field solution [Eq. (15)] is marked with the dashed line. (b) Trajectory-averaged steady-state half-chain entanglement entropy obtained with the two methods [see legend in (a)]. (c) Trajectory-averaged spin-wave density $\overline{\epsilon}$ in the steady state.
We also evaluate the performance of the spin-wave method on trajectory-averaged quantities. Fig. 3 (a) shows the time-evolution of the spin expectations $\overline{\langle\hat{S}^{x,y,z}\rangle}$ obtained by averaging spin-wave quantum trajectories. These are then benchmarked against the exact solution of the Lindblad master equation (14). As expected from the single-trajectory performance, the average dynamics is accurately reproduced. Notably, the decaying oscillations of the magnetization are due to the finite size effect and therefore require corrections beyond the mean-field level to capture. It is interesting to note that this decaying behavior is not displayed on the level of single trajectories, such as the one shown in Fig. 2 (a). This is typically observed in dissipative time crystals [114, 115, 116], and can be understood as the continuous monitoring preserving the coherence of the state. However, the different trajectories will have random relative phase shifts between each other which grow with time, such that at later times the different trajectories are no longer in phase, resulting in the decaying oscillations in the averaged quantities. This also implies that the averaged state should be highly mixed, where the different components of the mixture are states close to spin coherent states pointing in different directions. The resulting mixed state is hence far from a state that can be well approximated by the lowest-order Holstein-Primakoff expansion and the Gaussian approximation as the length of the averaged collective spin vector $\overline{\langle\hat{S}^{\alpha}\rangle}$ is only a fraction of the maximal value $S$. As a consequence, a deterministic version of the spin-wave theory with the same approximations applied to the averaged state, e.g., the one proposed in Ref. [97], will completely fail to capture the dynamics in this regime, cf. Appendix E for a comparison between our results and that given by the method in Ref. [97]. In Fig. 3 (b), we compare the trajectory-averaged entanglement entropy with that calculated from exact trajectories. Again, the two solutions are in good agreement, while the entanglement given by the exact simulation is slightly higher than the spin-wave results. This discrepancy is more noticeable in the presence of strong entanglement, as shown in Fig. 4, where we compare the steady-state solutions of the spin-wave method against the exact ones for a wide range of the drive $\omega$. Fig. 4 (c) shows the trajectory-averaged steady-state spin-wave density $\overline{\epsilon}$ associated with the spin-wave solutions. The peak in $\overline{\epsilon}$ corresponds to the maximum in the entanglement entropy [in panel (b)], where the difference between the spin-wave and exact solutions is also more pronounced. This illustrates the significance of the control parameter $\overline{\epsilon}$ as a signal for the validity of the spin-wave theory. On the other hand, the qualitative behavior of the entanglement is correctly captured despite the relatively high spin-wave density at its peak, and the magnetization predicted by the spin-wave method [panel (a)] remains accurate for all values of $\omega$ considered.
Figure 5: Results for the power-law interacting model in the long-range regime ($\alpha=0.2$, $J=0.1\kappa$) of different system sizes (colorbar shared across all panels). The dashed line marks the mean-field solution. (a) Steady-state average $z$ magnetization as a function of the drive $\omega$. (b) Steady-state of the trajectory-averaged half-chain entanglement entropy as a function of $\omega$. The vertical dotted lines mark the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 1.077\kappa$. Inset: scaling of the maximum entropy versus $N$ in linear-log scale. (c) Steady-state spin-wave density in log-linear scale. (d) Dynamics of the trajectory-averaged half-chain entanglement entropy for a driving value $\omega=1.06\kappa$, in linear-log scale.
Figure 6: Results for the power-law interacting model in the short-range regime ($\alpha=1.0$, $J=0.1\kappa$) of different system sizes (colorbar shared across all panels). (a) Steady-state average $z$ magnetization as a function of the drive $\omega$. The dashed line marks the mean-field solution. (b) Steady-state of the trajectory-averaged half-chain entanglement entropy as a function of $\omega$. The vertical dotted lines mark the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 1.077\kappa$. Inset: scaling of the maximum entropy versus $N$ in linear scale. (c) Steady-state spin-wave density in log-linear scale. (d) Dynamics of the trajectory-averaged half-chain entanglement entropy for a driving value $\omega=1.06\kappa$, in linear-log scale.
IV.2 Long-range case: $\alpha\neq 0$
To investigate the effect of a finitely long-range interaction and the entanglement dynamics in this regime, we apply the spin-wave method to the model with finite values of $\alpha$. We first consider a long-range model with $\alpha=0.2$. Fig. 5 (a) shows the steady-state expectation of the collective magnetization $\overline{\langle\hat{S}^{z}\rangle}$ as a function of the drive amplitude $\omega$ and the system size $N$. As we approach the thermodynamic limit by increasing $N$, the spin-wave solution converges towards the mean-field prediction and a continuous transition emerges separating the normal phase with $\overline{\langle\hat{S}^{z}\rangle}\neq 0$ and the time-crystal phase with $\overline{\langle\hat{S}^{z}\rangle}=0$. Fig. 5 (b) shows the behavior of the long-time-averaged half-chain entanglement entropy across this dissipative phase transition. The entanglement develops a logarithmic divergence as a function of $N$ at the critical point (as implied by the finite-size scaling analysis of the maximum value of the entanglement entropy), while it appears to be $N$-independent for drive values deep inside both phases. This suggests the emergence of an entanglement phase transition separating two area-law phases. This anomaly in entanglement also corresponds to a peak in the spin-wave density, as shown in Fig. 5 (c). When the system size $N$ increases, the spin-wave density is suppressed (including its peak value), which is a signature of the long-range nature of the model: in the $N\to\infty$ limit, a vanishing spin-wave density suggests that the system becomes equivalent to a mean-field (infinite-range) one. In Fig. 5 (d), we study the time evolution of the trajectory-averaged entanglement entropy at a driving value $\omega=1.06\kappa$, which is close to where the maximum steady-state entanglement is achieved for the finite system sizes we studied. Note that we are increasing $N$ exponentially between the different considered values, while the entanglement quickly reaches the steady-state value in all cases.
To study the effect of the interaction range on the entanglement dynamics, we also consider the short(er)-range case with $\alpha=1$ for comparison. In Fig. 6 we present the same physical quantities as considered in Fig. 5 for the long-range case, whereas here we are increasing $N$ in a linear scale between different system sizes considered. The steady-state magnetization $\overline{\langle\hat{S}^{z}\rangle}$, as shown in Fig. 6 (a), exhibits qualitatively similar behavior to the long-range case, while it shows more deviation from the mean-field prediction. This could result from more significant finite-size effects due to short-range interactions. (Note that the mean-field theory does not account for the interaction range parameter $\alpha$ and therefore gives identical predictions regardless of $\alpha$.) In sharp contrast, Fig. 6 (b) shows that the entanglement entropy of the short-range systems grows much faster with the system size $N$. Finite-size scaling suggests a volume law for the maximum entanglement entropy as well as in a vicinity around the critical point. The spin-wave density, as shown in Fig. 6 (c), also presents a qualitatively different behavior from the long-range case. Close to the maximal entanglement, the spin-wave density does not appear to decrease with the system size $N$, and remains around $\overline{\epsilon}\lesssim 0.2$. This suggests that the short-range model does not reduce to a mean-field (infinite-range) one even in the thermodynamic limit of $N\to\infty$, and that higher-order corrections to the theory are expected to have a more significant contribution. Finally, we show in Fig. 6 (d) the time dynamics of the trajectory-averaged entanglement entropy at $\omega=1.06\kappa$. Contrary to the fast saturation of entanglement in the long-range (or infinite-range) regime, the time it takes to reach the steady-state value scales rapidly with $N$. Due to the relatively high spin-wave density at the considered driving, we expect our result to hold only qualitatively.
The results presented for both the long- and short-range cases are in qualitative agreement with the observations made in [91] although the latter are for much more modest system sizes far from the thermodynamic limit. Here, the reported results obtained with the spin-wave method are up to $N=512$ long-range interacting spins, which is well beyond the exactly-solvable regime. We have therefore provided much stronger evidence to support the hypothesis that entanglement phase transition persists in the considered system and exhibits fast-saturating dynamics in the presence of sufficiently long-range interactions. On the other hand, our results also suggest the breakdown of this fast saturating property in the short-range regime.
V Experimental observability of monitored phases boosted by spin-wave quantum trajectories
The stochastic dynamics governed by Eq. (1) is continuous in time, while the experimental characterization of quantum trajectories typically requires a finite temporal binning, which effectively discretizes the evolution into small time steps. Therefore, we first discuss the discretized version of the dynamics in Eq. (1) and recast it in terms of the system and ancilla degrees of freedom, a description that is particularly tailored for digital quantum simulators. Afterwards, we discuss how the SWQT allows us to substantially mitigate the post-selection within the range of applicability of the semi-classical approximations.
V.1 From continuous to discrete monitored dynamics
@@
In view of the experimental implementations, for example on digital quantum simulators with discrete-time dynamics, we discuss in this subsection the discretized version of the dynamics in Eq. (1) by fixing $\Delta t$ a (small) time scale of the unitary and measurement steps, which is equivalent to the continuous-time case in the limit of $\Delta t\to 0$ (see also Refs. [118, 119, 25]). As shown in Appendix A, the stochastic master equation (1) can be cast in a diagonal form such that the noises ${\rm{d}}Z_{j}$ associated with different Lindblad jump operators are independent. We consider therefore the simplified problem with a single Lindblad jump operator $\hat{L}$, and focus on the dissipative part of the dynamics (fixing $\hat{H}=0$), namely
@@
Figure 7: Benchmark of the binary approximation of the Gaussian noise in the heterodyne-detection unraveling. We compare trajectory-averaged quantities obtained from different methods and different types of noise (see legend), namely 1) the spin-wave quantum trajectory (SWQT) method with binarized noise, 2) the exact solution of the stochastic master equation with Gaussian noise and 3) the exact solution but with binary noise replacing the gaussian noise. Panel (a) compares the results on a linear quantity $\overline{\langle\hat{S}^{z}\rangle}$. Panel (b) and (c) show the benchmark on two different nonlinear quantities, the half-system entanglement entropy $\overline{S_{E}}$ and the squared magnetization $\overline{\langle\hat{S}^{z}\rangle^{2}}$ respectively. Parameters: $\omega=1.3\kappa$, $J=0.1\kappa$, $\alpha=0$ and $S=64$.
@@
In a similar fashion, we have $\overline{\Delta R_{x}^{2}}=\overline{\Delta R_{y}^{2}}=\Delta t$ and $\overline{\Delta R_{x}\Delta R_{y}}=0$. Putting all together, and defining the zero-mean stochastic processes ${\rm{d}}X\equiv\Delta R_{x}-\overline{\Delta R_{x}}$ and ${\rm{d}}Y=\Delta R_{y}-\overline{\Delta R_{y}}$, which satisfy $\overline{{\rm{d}}X^{2}}=\overline{{\rm{d}}Y^{2}}=\Delta t$ and $\overline{{\rm{d}}X{\rm{d}}Y}=0$, we obtain the final expression after simple algebra  [^3] Note that since the random variable $\Delta R_{x}\Delta R_{y}$ has neglible variance (of the order $\Delta t^{2}$), we replace it with the mean: $\Delta R_{x}\Delta R_{y}=\overline{\Delta R_{x}\Delta R_{y}}=0$ [123]. Therefore, terms proportional to $\Delta R_{x}\Delta R_{y}$ are dropped in this derivation.,
@@
which is equivalent to Eq. (47) when substituting ${\rm{d}}Z=({\rm{d}}X+{\rm{i}}{\rm{d}}Y)/\sqrt{2}$.
We verify the validity of this binary approximation of the Gaussian noise by computing the trajectory-averaged quantities including both linear and non-linear functions of the state and compared the results with the exact solution obtained with the Gaussian noise. An example of the benchmark is shown in Fig. 7, where we computed the magnetization $\overline{\langle\hat{S}^{z}\rangle}$, the half-system entanglement entropy $\overline{S_{E}}$ and the nonlinear quantity $\overline{\langle\hat{S}^{z}\rangle^{2}}$ from three different methods: 1) Spin-wave quantum trajectories with binarized noise, 2) exact solution of the stochastic master equation with Gaussian noise and 3) exact solution with the binary approximation of the noise. In all cases, we observe excellent agreement among the results obtained in these different ways.

### v2

@@ IV.1.1 Infinitesimal increments
The stochastic master equation (1) presented in Sec. II.1 determines the time evolution of the expectation value of operators. For a time-independent operator $\hat{O}$, the expectation $\langle\hat{O}\rangle={\rm{Tr}}[\hat{\rho}\hat{O}]$ evolves as follows,
@@
In the equations above, the Hamiltonian $\hat{H}$ and the dissipation operators $\hat{L}_{i}$ should be expressed in terms of the bosonic operators $\hat{b}_{i}$ using the substitution rules defined in Eq. (9). The Gaussian approximation then allows evaluating the expectation value of every term beyond quadratic order in $\hat{b}_{i}$ in terms of one- and two-point correlators, i.e. $\beta_{i}$, $u_{ij}$ and $v_{ij}$, thanks to the Wick theorem. The increments $\delta\beta_{i}$, $\delta u_{ij}$, and $\delta v_{ij}$ are then calculated using the discretized versions of the equations above. This can be achieved with the substitution ${\rm{d}}t\to\delta t$ for a sufficiently small time step $\delta t$. The noise is approximated with ${\rm{d}}\mbox{w}_{i}\to\delta\mbox{w}_{i}=\sqrt{\delta t}(X_{i}+{\rm{i}}Y_{i})$, where the pair of random vectors $(\mathbf{X},\mathbf{Y})$ at every time step is drawn from a multivariate real Gaussian distribution with zero mean and the following covariance matrix,
@@ IV.1.2 Re-alignment of the frame
as a direct implication of Eq. (9). This can be achieved by moving the frame smoothly along a path $\theta(\tau),\phi(\tau)$ parametrized by some parameter $\tau$. The unitary transformation $\hat{U}(\theta,\phi)$ therefore becomes $\tau$-dependent,
@@
This equation can be integrated analytically by considering the path on the Bloch sphere from $(\theta_{1},\phi_{1})$ to $(\theta_{2},\phi_{2})$ following the two segments [as illustrated in Fig. 1 (c)]:
@@
The solution is then
$\hat{b}_{i}(\theta_{2},\phi_{2})=$ $\leavevmode\nobreak\ \mathopen{}\mathclose{{}\left[\sqrt{\dfrac{s}{2}}\tan\theta_{1}+\hat{b}_{i}(\theta_{1},\phi_{1})}\right]{\rm{e}}^{-{\rm{i}}\Delta\phi\cos\theta_{1}}$ (26)
@@
where $\Delta\theta\equiv\theta_{2}-\theta_{1}$ and $\Delta\phi\equiv\phi_{2}-\phi_{1}$. This equation immediately implies the evolution of $\underline{\beta}$, upon taking the expectation of both sides,
@@
With the angular increments in hand, the first moments $\beta_{i}$ of the Gaussian ansatz can be updated directly using Eq. (26) with the operator $\hat{b}_{i}$ replaced by its expectation value $\beta_{i}$. This equation also implies the evolution of the fluctuation operators $\hat{\delta}_{i}=\hat{b}_{i}-\beta_{i}$, which is simply
@@
IV.2 From continuous to discrete monitored dynamics
@@
We discuss in this section the discretized version of the dynamics in Eq. (1) by fixing $\Delta t$ a (small) time scale of the unitary and measurement steps, which is equivalent to the continuous-time case in the limit of $\Delta t\to 0$ (see also Refs. [72, 73, 10]). As shown in Supplementary Note A, the stochastic master equation (1) can be cast in a diagonal form such that the noises ${\rm{d}}Z_{j}$ associated with different Lindblad jump operators are independent. We consider therefore the simplified problem with a single Lindblad jump operator $\hat{L}$, and focus on the dissipative part of the dynamics (fixing $\hat{H}=0$), namely
@@
In a similar fashion, we have $\overline{\Delta R_{x}^{2}}=\overline{\Delta R_{y}^{2}}=\Delta t$ and $\overline{\Delta R_{x}\Delta R_{y}}=0$. Putting all together, and defining the zero-mean stochastic processes ${\rm{d}}X\equiv\Delta R_{x}-\overline{\Delta R_{x}}$ and ${\rm{d}}Y=\Delta R_{y}-\overline{\Delta R_{y}}$, which satisfy $\overline{{\rm{d}}X^{2}}=\overline{{\rm{d}}Y^{2}}=\Delta t$ and $\overline{{\rm{d}}X{\rm{d}}Y}=0$, we obtain the final expression after simple algebra,
@@
which is equivalent to Eq. (32) when substituting ${\rm{d}}Z=({\rm{d}}X+{\rm{i}}{\rm{d}}Y)/\sqrt{2}$. We benchmark the validity of this binary approximation in Supplementary Note D.6.